OpenStudy (anonymous):

please help this is my first post The perimeter of a rectangle is 50 cm and the length is 4 cm longer than the width. Write and solve the system of equations using the substitution method. A. width = 10 cm, length= 14 cm B. width = 10 1/2 cm, length= 14 1/2 cm C. width = 11 cm, length= 15 cm D. width = 23 cm, length= 27 cm

Vocaloid (vocaloid):

let W = width L = length perimeter = 2(L+W) L = W + 4 do you think you can take it from here?

OpenStudy (anonymous):

no i dont understand any of this lol thats why im a senior retaking algebra one

Vocaloid (vocaloid):

well, to find the perimeter of a rectangle, we add the width and length, then multiply everything by 2, so perimeter = 2(L+W) make sense so far?

OpenStudy (anonymous):

yes that makes sense

Vocaloid (vocaloid):

ok, the problem also tells us that "length is 4 more than the width" so L = W + 4, make sense so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

great, now let's put them together and go from there perimeter = 50 = 2(L+W) L = W + 4 now, the second equation tells us that we can replace "L" with "W+4" whenever we see L, so we can plug that into the first equation, giving us: 50 = 2(L+W) = 2(W + 4 + W) with me so far?

OpenStudy (anonymous):

kinda

Vocaloid (vocaloid):

well, the second equation tells us that "length" is the same as "width+4" so we can replace "length" with "width + 4". the reason why we do this is to make all the letters in the first equation the same, so we can find width 50 = 2(W + 4 + W) we can put W + W together to get 2W, which gives us 5= 2(2W+4), with me so far?

OpenStudy (anonymous):

yes im with you

Vocaloid (vocaloid):

ok, so we have 50 = 2(2W+4) we can divide both sides by 2 to get 25 = (2W+4) do you think you can solve for W now?

OpenStudy (anonymous):

i think i got it

Vocaloid (vocaloid):

ok, give it a shot and tell me what W is I'll guide you if you get stuck

OpenStudy (anonymous):

ok i have no idea

Vocaloid (vocaloid):

no worries 25 = (2W+4) subtract 4 from each side, what do you get?

OpenStudy (anonymous):

wait is the answer D

Vocaloid (vocaloid):

nope, not D, please try the problem with me 25 = (2W+4) subtract 4 from both sides of the equals sign

OpenStudy (anonymous):

subtract 4 from 25 thats 21 ok

OpenStudy (anonymous):

its B

Vocaloid (vocaloid):

what about the right side? 2W + 4 - 4 = ?

OpenStudy (anonymous):

its B :)

Vocaloid (vocaloid):

great

OpenStudy (anonymous):

thank you!!!!!!!!